OpenStudy (anonymous):

Help please! The number of visitors to certain web site triples every month. The number of visitors is modeled by the expressions 8100 * 3^m , where m is the number of months after the number of visitors was measured. Evaluate the expression m = -4. What does the value of the expression represent in the situation .

OpenStudy (campbell_st):

well m = -4 you have to remember negative powers indicate fractions \[x^{-a} = \frac{1}{x^a}\] apply that index law here and you get \[8100 \times 3^{-4} = \frac{8100}{3^4}\] all you need to do is evaluate hope it helps

OpenStudy (anonymous):

I evaluate 8100/3^4 .. ? @campbell_st

OpenStudy (campbell_st):

well what is \[3^4 = 3 \times 3 \times 3 \times 3 = ?\]